OpenStudy (anonymous):

Need Help please!!! What's the answer to x(y+2)=xy+xz

OpenStudy (anonymous):

Both Please

OpenStudy (anonymous):

x(y+2) = xy + xz xy +2x = xy+xz subtract xy from both sides assuming xy isnt 0 2x = xz divide by x (assuming x isnt 0) z=2

OpenStudy (anonymous):

Thanks